VT R Efficiency Pin loss Pin-P R1 Key A R2 V1 ::::::::::::::::::31 i2v:::::::::::: R4 2kQR R3 GND :::::::::::s2::::::::::::::::32RO: Key A ::::::::::::::::E: Fig-1 RT
media%2Fc5a%2Fc5a59e67-c8cf-4eff-8fd1-49
Answer #1

a) i)

When switch 1 and 2 are both selected, it completes the circuit,

So the effective resistance would be,

R1 and R3 in series,

So R = 1+2= 3 ohm.

ii) When switch 2 is not selected, the entire circuit becomes open.

hence effective resistance = 0

b) Power in the circuit,

when switch 1 is selected is,

P = \frac{V^{2}}{R}

P = \frac{12^{2}}{3} = 48 watt

when switch 2 is selected,

P =1/0 = infinite.